Sub 24
Ques. 28

The lines with equations x = 2, x = 5, y = 7, and y = 3 form a rectangle. What is the area of this rectangle?
1
Expert's answer
2012-05-15T10:48:33-0400
delta y=7-4=4
delta x=5-2=3
So the area is 3*4=12